# Seungwon Baek b,c    P. Ko d    Yuji Omura e    and Chaehyun Yu

### v1

@@ 4.1 Parameters
$1\leq\tan\beta\leq 100\quad\textrm{and}\quad 1\leq v_{\Phi}/\textrm{GeV}\leq 10^{4}.$ (35)
However, it turns out that $1\leq\tan\beta\leq 10$ and $5\times 10^{3}\leq v_{\Phi}/$GeV $\leq 10^{4}$ are sufficient for the scanning of the parameters. The ranges of the masses of the extra scalar bosons are taken to be $10\leq m_{H,A,H^{\pm}}/$GeV $\leq 1000$, respectively. The Yukawa couplings of the extra quark and lepton are in the ranges of $0\leq y_{D,L}\leq 4\pi$, respectively, but for simplicity we set $y_{D}=y_{L}=1$. We note that essential features from the analysis are not modified even though larger Yukawa couplings are chosen. The $U(1)_{H}$ gauge coupling must not be large because of constraints from the $Z^{\prime}$ boson search at the LHC. We take $g_{X}$ to be in the range of $0.01\leq g_{X}\leq 0.1$. Then, the $Z^{\prime}$ mass, $m_{Z^{\prime}}$, is written in terms of the parameters
@@ 4.2 Analysis
In Fig. 1, we present the signal strengths $\mu_{\gamma\gamma}$ and $\mu_{\tau\tau}$ in both Type-II models with extra fermion contributions in the loop (left) and without those contributions (right), respectively. The extra fermion contribution in the loop can especially modify branching ratios of $\tilde{h}\to gg$ or $\gamma\gamma$. Thus, the production or decay rates of $\tilde{h}$ can be affected by the extra fermion loop contributions. The two dashed lines denote $1\sigma$ and $2\sigma$ regions for the combined experimental values in ATLAS and CMS, respectively. We find that the expectation in our Type-II models is inconsistent with the experiments within $1\sigma$, while some points are consistent within $2\sigma$. In particular, it is difficult to enhance the $\mu_{\tau\tau}$ because of effective Yukawa couplings of the 96 GeV scalar boson to the top quarks and tau lepton in the Type-II models:
@@
Because the two effective Yukawa couplings are the same, both production cross sections for $gg\to\tilde{h}\to\gamma\gamma$ and $gg\to\tilde{h}\to\tau\tau$ tend to increase simultaneously as $\sin\alpha_{1}$ increases. It turns out that the branching ratio for $\tilde{h}\to\tau\tau$ can reach about $0.5$ in the allowed parameter space. Then, $\mu_{\tau\tau}$ can be enhanced compared to the Type-II models and some parameter sets can be consistent with the experimental values within $1\sigma$.
Figure 3: Signal strengths $\mu_{\gamma\gamma}$ and $\mu_{\tau\tau}$ in the Type-Y models. The blue points correspond to the model including the extra fermions in the loop, while the red ones to the model without their contributions.
In order to find effects of the extra fermions in the loop, we present the signal strengths $\mu_{\gamma\gamma}$ and $\mu_{\tau\tau}$ in the Type-Y models with extra fermion-contributions in the loop (blue) and without those contributions (red) in Fig. 3, respectively. Each point corresponds to its own parameter set. Thus, overlapped points means that the contributions from the extra fermions may not be crucial. However, it is apparent that some of blue points are not overlapped with red points and vice versa. The former indicates that the parameter set is consistent with all the experiments, but if the contributions from the extra fermions are ignored, it cannot satisfy one of constraints. The red points which do not overlap with the blue points mean that the parameter set seems to satisfy all constraints without extra fermions. However, if the extra fermion contributions are included, they cannot be candidates for the resonance signals at 96 GeV.
Figure 4: Signal strengths $\mu_{bb}$ and $\mu_{\tau\tau}$ in the Type-II model (left) and in the Type-Y model (right) with extra fermion contributions in the loop.
In Fig. 4, we present the signal strengths $\mu_{bb}$ and $\mu_{\tau\tau}$ in the Type-II model (left) and in the Type-Y model (right) with extra fermion contributions in the loop. The effective couplings of $\tilde{h}$ to the bottom quark are
@@
In the Type-II 2HDM, the effective Yukawa coupling, $g_{\tilde{h}}^{b,\textrm{II}}$, of $\tilde{h}$ to the bottom quark is the same as that to the tau lepton. Thus, $\mu_{bb}$ tends to be proportional to $\mu_{\tau\tau}$ as shown in the left figure of Fig. 4. It is obvious that $\mu_{bb}$ in the Type-II models is preferred to be similarly small like $\mu_{\tau\tau}$ in the allowed parameter space. On the other hand, in the Type-Y models, the effective Yukawa coupling, $g_{\tilde{h}}^{b,\textrm{Y}}$, of $\tilde{h}$ to the bottom quark is different from that to the tau quark. As shown in the right figure of Fig. 4, a wider region is allowed in the Type-Y models and some parameters could be within $1\sigma$ from the central value of the experiments.
Figure 5: Signal strengths $\mu_{\gamma\gamma}$ and $\mu_{\tau\tau}$ in the Type-II model (left) and in the Type-Y model (right), respectively. The blue points are the same as those in the left figure of Figs. 1 and 2. The red points satisfy another constraint from $gg\to tt\tau\tau$ at ATLAS.
4.3 Constraints from $t\bar{t}\tau^{+}\tau^{-}$ production at the LHC
The 96 GeV scalar boson couples to both the top quark and tau lepton. Thus the production of $t\bar{t}\tau^{+}\tau^{-}$ at the LHC can constrain the scalar boson Iguro:2022dok  [^1] See also Ref. Iguro:2022fel .. Based on the ATLAS Run 2 full data ATLAS:2022yrq , the authors in Ref. Iguro:2022dok have proposed that the Yukawa couplings, $\rho_{tt,\tau\tau}^{\tilde{h}}$, normalized by $\sqrt{2}$, should satisfy either
$\rho_{\tilde{h}}^{tt}$ $=$ $\frac{g_{\tilde{h}}^{t,\textrm{Y}}m_{t}}{v}<0.2,$ (41)
or
$\rho_{\tilde{h}}^{\tau\tau}$ $=$ $\frac{g_{\tilde{h}}^{\tau,\textrm{Y}}m_{\tau}}{v}<0.002,$ (42)
respectively. It is worthwhile to mention that these constraints are based on the analysis of the tree-level calculation for the $t\bar{t}\tau^{+}\tau^{-}$ production Iguro:2022dok . With QCD corrections to the process, the bounds may be slightly modified.
In Fig. 5, we present the signal strengths $\mu_{\gamma\gamma}$ and $\mu_{\tau\tau}$ in the Type-II model (left) and in the Type-Y model (right) with extra fermion contributions, respectively. The blue points are the same as as those in the left figure of Figs. 1 and 2. The red points satisfy constraints from $gg\to t\bar{t}\tau^{+}\tau^{-}$ at ATLAS in Eqs. (41) and (42). In the Type-II case, we find that the constraints from $gg\to t\bar{t}\tau^{+}\tau^{-}$ do not provide any significant bounds to the model. However, in the Type-Y case, the constraints are very stringent and most of allowed spaces are rejected. It turns out that the allowed region in the Type-Y model is about $1.3\sigma$ as much as possible. In the Type-II model, the allowed region is about $1.5\sigma$, which would be comparable to that in the Type-Y model. Therefore, we find that the Type-Y model is slightly better than the Type-II model to explain the 96 GeV resonances. However, both models are consistent with the resonances within similar statistical errors if the constraints from $gg\to t\bar{t}\tau^{+}\tau^{-}$ at ATLAS are considered.
@@ 4.4 Benchmark points
We have scanned the parameters, but we see that the allowed parameter range is limited because of the strong constraints, as discussed in Sec. 4.1. In Table 5, we introduce some benchmark points that explicitly show our predictions. Those parameter sets evade the bound from $t\bar{t}\tau^{+}\tau^{-}$, discussed in Sec. 4.3, and can enhance $\mu_{\gamma\gamma}$ and $\mu_{\tau\tau}$. $\mu_{bb}$ is within 2 $\sigma$ in the cases other than the point 3.
There is no big difference between in the Type-II 2HDM and in the Type-Y 2HDM. The constraints do not allow the mixing among scalars to be large. In fact, the mixing angles, $\alpha_{1}$, $\alpha_{2}$ and $\alpha_{3}$, in all cases, correspond to ${\cal O}(1)$ $|\sin\alpha_{1,2,3}|$. Then, $\tilde{h}$, whose mass is 96 GeV, is dominated by $h_{\Phi}$. The SM-like Higgs boson, $h$, dominantly consists of $h_{2}$. $\tan\beta$ is also small, so that $|\sin\beta|$ is also ${\cal O}(1)$ and the heavy scalars, $A$ and $H$, mainly consists of $\chi_{1}$ and $h_{1}$: $A\sim\chi_{1}$ and $H\sim h_{1}$. While $|\sin\alpha_{1,2,3}|$ and $|\sin\beta|$ are ${\cal O}(1)$, $|\cos\alpha_{1,2,3}|$ and $|\cos\beta|$ are at least ${\cal O}(0.1)$ to enhance $\mu_{\gamma\gamma}$ and $\mu_{\tau\tau}$. Thus, $\tilde{h}$ is approximately expressed as $\tilde{h}\sim h_{\Phi}+{\cal O}(0.1)\,h_{2}$ in the benchmark points. As shown in Table 5, $g_{X}$ is very small to evade too large deviation of the $\rho$ parameter, so the mixing parameters are relevant to the signal strengths. In both cases, $\mu_{\gamma\gamma}$ and $\mu_{\tau\tau}$ can be within 2 $\sigma$ as shown in Fig. 5, although $\mu_{\tau\tau}$ in the Type-II 2HDM is smaller than that in the Type-Y 2HDM.
@@
$\alpha_{1}$ 1.35 1.37 1.53 1.46
$\alpha_{2}$ 1.21 1.23 1.34 1.29
$\alpha_{3}$ 4.67 4.65 4.22 4.54
$\tan\beta$ 3.20 3.48 1.60 3.28
$v_{\Phi}$ (GeV) $7350.70$ 7309.07 5426.80 8894.68
$m_{H}$ (GeV) 649.56 707.46 645.07 586.47
$m_{A}$ (GeV) 631.48 694.74 503.50 552.09
$m_{H_{\pm}}$ (GeV) 649.31 672.13 605.20 601.70
@@
$g_{X}$ $4.76\times 10^{-2}$ $6.11\times 10^{-2}$ $4.51\times 10^{-2}$ $5.65\times 10^{-2}$
$M_{Z^{\prime}}$ (GeV) 350.16 446.83 245.08 502.82
$\mu_{bb}$ $0.11$ $9.75\times 10^{-2}$ $1.02\times 10^{-3}$ $2.83\times 10^{-2}$
$\mu_{\gamma\gamma}$ 0.24 0.23 0.20 0.30
$\mu_{\tau\tau}$ 0.14 0.13 0.50 0.34
@@ 5 Summary
We studied the di-photon and di-$\tau$ channels at the LHC as well as the $b\bar{b}$ channel at LEP, in the Type-II and Type-Y 2HDMs with gauged $U(1)_{H}$ symmetry. The $U(1)_{H}$ gauge symmetry distinguishes one Higgs doublet from the other, and naturally suppresses tree-level FCNCs. One extra scalar charged under $U(1)_{H}$ is also introduced, so that there are three CP-even neutral scalars at low energy. In our study, the mass of lightest neutral scalar, denoted by $\tilde{h}$, was fixed at 96 GeV that corresponds to the mass region where the mild excesses are reported by the ATLAS and CMS collaborations. In our models, extra fermions as well as extra scalars contribute to the di-photon and di-$\tau$ channels. The extra gauge boson deviates $\rho$ parameter from $1$. We surveyed our predictions of the signal strengths, $\mu_{bb}$, $\mu_{\gamma\gamma}$ and $\mu_{\tau\tau}$, scanning parameters. The $Z^{\prime}$ couplings are strongly constrained by the $\rho$ parameter and the $Z^{\prime}$ searches, while the contribution of extra fermions affects the signal strengths in some parameter sets. We saw that the dominant contributions to $\mu_{\gamma\gamma}$ and $\mu_{\tau\tau}$ are from the mixing among three neutral scalars, not from the extra new fermions. As discussed in Sec. 4, the lightest neutral scalar is mainly composed of the scalar from $\Phi$. This result is consistent with the previous work Biekotter:2023oen . Our work reveals the underlying theory of the 2HDM with a complex scalar field Biekotter:2023oen , and explicitly shows how large contributions of the extra fields predicted by extending the symmetry, that distinguishes two Higgs fields, to gauged $U(1)_{H}$ symmetry can be. The predictions of the signal strengths, in fact, deviate from them of the 2HDMs without extra fermions. Further study would be required to reveal which setup is preferred by the excesses around 96 GeV.

### v2

@@ 4.1 Parameters
$1\leq\tan\beta\leq 30\quad\textrm{and}\quad 5\times 10^{3}\leq v_{\Phi}/\textrm{GeV}\leq 10^{4}.$ (36)
The ranges of the masses of the extra scalar bosons are taken to be $500\leq m_{H,A,H^{\pm}}/$GeV $\leq 1500$, respectively, considering the bounds from flavor physics. We have checked broader ranges for $\tan\beta$, $v_{\Phi}$, and the scalar boson masses, but we have found that it does not change the overall behavior of signal strengths in allowed regions. We have also limited the angles in the ranges of $0.5\leq\alpha_{1}\leq 1.5$, $1\leq\alpha_{2}\leq 2$, and $1\leq\alpha_{3}\leq 2$ because no allowed regions exist outside these ranges in the first scan with the entire range, or the other allowed regions does not change the overall feature. At first, the Yukawa couplings of the extra quark and lepton are in the ranges of $0\leq y_{D,L}\leq 4\pi$, respectively, but for simplicity we set $y_{D}=y_{L}=1$. We note that essential features from the analysis are not modified even though larger Yukawa couplings are chosen. The $U(1)_{H}$ gauge coupling must not be large because of constraints from the $Z^{\prime}$ boson search at the LHC. We take $g_{X}$ to be in the range of $0.01\leq g_{X}\leq 0.1$. Then, the $Z^{\prime}$ mass, $m_{Z^{\prime}}$, is written in terms of the parameters
@@ 4.2 Analysis
In Fig. 1, we present the signal strengths $\mu_{\gamma\gamma}$ and $\mu_{\tau\tau}$ in both Type-II models with extra fermion contributions in the loop (left) and without those contributions (right), respectively. The extra fermion contribution in the loop can especially modify branching ratios of $\tilde{h}\to gg$ or $\gamma\gamma$. Thus, the production or decay rates of $\tilde{h}$ can be affected by the extra fermion loop contributions. The two dashed lines denote $1\sigma$ and $2\sigma$ regions for the combined experimental values in ATLAS and CMS, respectively. The blue and red points represent parameter sets consistent with the constraints discussed in the previous section. The only distinction between these two sets lies in the flavor physics constraints applied. The blue points satisfy the HFLAV cut for the $B\to X_{s}\gamma$ decays, which imposes a lower bound of $m_{H^{\pm}}\gtrsim 500$ GeV. Since the $B_{s}\to\mu^{+}\mu^{-}$ constraint is more stringent than those from other observables in both low and high $\tan\beta$ regimes Li:2024kpd , we impose the bound provided in Ref. Li:2024kpd . The red points satisfy the tight cut, which is based on the constraints in Ref. Li:2024kpd . We note that the lower limit of $m_{H^{\pm}}$ in this cut is about 800 GeV while the $B_{s}\to\mu^{+}\mu^{-}$ constraint is the same as those in the HFLAV cut. Since the tight cut is more stringent than the HFLAV cut, all the red points overlap the blue points. The yellow points represent parameter sets that satisfy all constraints except for those from HiggsSignals, while adhering to the HFLAV cut for the flavor physics constraints. We note that the constraints from $t\bar{t}\tau^{+}\tau^{-}$ production at the LHC are applied to both red and blue points, while they are not to the yellow points.
We find that the expectation in our Type-II models is inconsistent with the experimental results within $1\sigma$, while some regions are consistent within $2\sigma$. In particular, it is difficult to enhance the $\mu_{\tau\tau}$ because of effective Yukawa couplings of the 96 GeV scalar boson to the top quarks and tau lepton in the Type-II models:
@@
As we see in Fig. 1, the region allowed by the HFLAV cut is broader than that consistent with the tight cut, in the Type-II models. However, the total signal strengths do not differ significantly in the two cases. We find that the constraint from HiggsSignals is very stringent, excluding substantial portion of the parameter space. Without these constraints, the model prediction could deviate by up to approximately $1.3\sigma$ from the experimental central values. This situation is similar in both models, with and without the extra fermion contributions in the loop. Despite this, it is apparent that the signal strengths are distinctive in both models. This implies that a parameter set consistent within $2\sigma$ with the extra fermion contributions might be excluded if the extra fermions are ignored in the loop, and vice versa. Thus, the extra fermions may play a role in the search for the 96 GeV boson.
@@
Since the two effective Yukawa couplings are identical, both production cross sections for $gg\to\tilde{h}\to\gamma\gamma$ and $gg\to\tilde{h}\to\tau\tau$ tend to increase simultaneously as $\sin\alpha_{1}$ increases. This behavior is indeed observed for both the blue and red points, which adhere to the HFLAV and tight cuts, respectively. Moreover, compared to the Type-II models, the $\mu_{\tau\tau}$ signal strength can be enhanced and the deviation from the experimental central values can fall within the $1\sigma$ range under the HFLAV cut. Even under the tight cut, the deviation can reach up to approximately $1.2\sigma$ from the central values. As shown in Fig. 2, the effects of extra fermions in the loop are very restrictive, but one can observe a small difference between two models with and without extra fermions in the loop. We find that some points which are excluded the model without extra fermions could be consistent within $2\sigma$ by including the extra fermions in the loop, similar to the Type-II models.
Figure 3: Signal strengths $\mu_{bb}$ and $\mu_{\tau\tau}$ in the Type-II model with extra fermions in the loop (left) and without them (right).
Figure 4: Signal strengths $\mu_{bb}$ and $\mu_{\tau\tau}$ in the Type-Y model with extra fermions in the loop (left) and without them (right).
In Figs. 3 and 4, we present the signal strengths $\mu_{bb}$ and $\mu_{\tau\tau}$ in the Type-II and Type-Y model 2HDMs, respectively. The extra fermion contributions are included in the left figures, while they are not included in the right figures. The effective couplings of $\tilde{h}$ to the bottom quark are
@@
In the Type-II 2HDM, the effective Yukawa coupling, $g_{\tilde{h}}^{b,\textrm{II}}$, of $\tilde{h}$ to the bottom quark is the same as that to the tau lepton. Thus, $\mu_{bb}$ tends to be proportional to $\mu_{\tau\tau}$ as shown in the left figure of Fig. 3. It is obvious that $\mu_{bb}$ in the Type-II models is preferred to be similarly small like $\mu_{\tau\tau}$ in the allowed parameter space. On the other hand, in the Type-Y models, the effective Yukawa coupling, $g_{\tilde{h}}^{b,\textrm{Y}}$, of $\tilde{h}$ to the bottom quark is different from that to the tau quark. As shown in Fig. 4, a wider region is allowed in the Type-Y models.
In our models, the extra fermions are required to ensure anomaly cancellation. These fermions can interact with both gauge bosons and scalar bosons. As a result, the branching ratios of scalar bosons into gauge bosons can be affected by extra fermions involving loop diagrams. However, as shown in the figures in this section, the overall features of the models remain unchanged by the inclusion of extra fermion contributions. Therefore, the impact of the extra fermions appears to be insignificant for the 96 GeV resonances. This is because the parameters in the scalar potential are tightly constrained by experimental data, such as flavor physics observables and precision measurements of the SM Higgs boson. The extra fermions play only a minor role in both types of experiments. However, we note that certain detailed features of the models may differ depending on whether the extra fermions are included. The signal strengths, $\mu_{\gamma\gamma}$ and $\mu_{\tau\tau}$ depend on the branching ratios of $\tilde{h}$ decays into $gg$ and/or $\gamma\gamma$. Therefore, the extra fermion contributions to these branching ratios may become significant if a given parameter point lies near the boundary of the $2\sigma$ allowed region.
@@ 4.3 Benchmark points
[…]
There is no big difference between in the Type-II 2HDM and in the Type-Y 2HDM, when charged Higgs is heavy, i.e., $m_{H^{\pm}}>800$ GeV. We note that the parameter values of point 1 in Table 5 are the same as them of point 3. We can see the difference of the mixing dependence on the $\tau$ couplings in the signal strengths. The mixing angles, $\alpha_{1}$, $\alpha_{2}$ and $\alpha_{3}$, in all cases, correspond to ${\cal O}(1)$ $|\sin\alpha_{1,2,3}|$. Then, $\tilde{h}$, whose mass is 96 GeV, is dominated by $h_{\Phi}$, while the SM-like Higgs boson, $h$, dominantly consists of $h_{1}$ and $h_{2}$. In all points in Table 5, $\tan\beta$ is small, and $|\cos\alpha_{1,2}|$ are larger than $0.1$ to enhance $\mu_{\gamma\gamma}$ and $\mu_{\tau\tau}$. As shown in Table 5, $g_{X}$ is very small to evade too large deviation of the $\rho$ parameter, so, the mixing parameters are relevant to the signal strengths.
@@
$\alpha_{1}$ 1.01 0.97 1.01 1.36
$\alpha_{2}$ 1.88 1.25 1.88 1.90
$\alpha_{3}$ 1.40 1.48 1.40 1.30
$\tan\beta$ 1.20 1.21 1.20 1.98
$v_{\Phi}$ (GeV) 7459.47 8646.47 7459.47 9899.79
$m_{H}$ (GeV) 900.36 749.49 900.36 652.26
$m_{A}$ (GeV) 713.55 630.49 713.55 549.93
$m_{H^{\pm}}$ (GeV) 877.49 692.21 877.49 605.38
@@
$g_{X}$ $3.83\times 10^{-2}$ $7.29\times 10^{-2}$ $3.83\times 10^{-2}$ $1.61\times 10^{-2}$
$\mu_{\gamma\gamma}$ 0.13 0.15 0.13 0.27
$\mu_{\tau\tau}$ $9.07\times 10^{-2}$ 0.13 0.15 0.35
$\mu_{bb}$ $8.59\times 10^{-2}$ $9.25\times 10^{-2}$ $8.08\times 10^{-2}$ $5.23\times 10^{-2}$
@@ 5 Summary
We studied the di-photon and di-$\tau$ channels at the LHC as well as the $b\bar{b}$ channel at LEP, in the Type-II and Type-Y 2HDMs with gauged $U(1)_{H}$ symmetry. We can simply add one extra Higgs doublet to the SM. Such a simple extension, however, causes FCNCs at the tree level, if two Higgs doublets are not distinguished. Moreover, extra scalars predicted by the two Higgs doublets usually reside around the EW scale, so that the extended SMs are easily excluded. In the conventional 2HDMs (e.g., the Type-II 2HDM and the Type-Y 2HDM), softly broken $Z_{2}$ symmetry is imposed on fields and the strong constraints are evaded. In our model, we see that softly broken $Z_{2}$ symmetry is originated from the spontaneous symmetry breaking of gauged $U(1)_{H}$ symmetry. In this sense, our model can be interpreted as the underlying theory of the conventional 2HDMs. Moreover, the 2HDM with a singlet scalar is also recently studied, motivated by the excesses around 96 GeV. Our 2HDM with gauged $U(1)_{H}$ symmetry also predicts a singlet scalar, so it is worthwhile to study the signals related to the resonance around 96 GeV in our models. In our model, the $U(1)_{H}$ gauge symmetry distinguishes one Higgs doublet from the other, and naturally suppresses tree-level FCNCs. One extra scalar charged under $U(1)_{H}$ is also introduced, so that there are three CP-even neutral scalars at low energy. In our study, the mass of lightest neutral scalar, denoted by $\tilde{h}$, was fixed at 96 GeV that corresponds to the mass region where the mild excesses are reported by the ATLAS and CMS collaborations. In our models, extra fermions as well as extra scalars contribute to the di-photon and di-$\tau$ channels. The extra gauge boson deviates $\rho$ parameter from $1$. We surveyed our predictions of the signal strengths, $\mu_{bb}$, $\mu_{\gamma\gamma}$ and $\mu_{\tau\tau}$, scanning parameters. The $Z^{\prime}$ couplings are strongly constrained by the $\rho$ parameter and the $Z^{\prime}$ searches, while the contribution of extra fermions affects the signal strengths in some parameter sets. We saw that the dominant contributions to $\mu_{\gamma\gamma}$ and $\mu_{\tau\tau}$ are from the mixing among three neutral scalars, not from the extra new fermions. As discussed in Sect. 4, the lightest neutral scalar is mainly composed of the scalar from $\Phi$. This result is consistent with the previous work Biekotter:2023oen . Our work reveals the underlying theory of the 2HDM with a complex scalar field Biekotter:2023oen , and explicitly shows how large contributions of the extra fields predicted by extending the symmetry, that distinguishes two Higgs fields, to gauged $U(1)_{H}$ symmetry can be. The predictions of the signal strengths, in fact, deviate from them of the 2HDMs without extra fermions. Further study would be required to reveal which setup is preferred by the excesses around 96 GeV. The improved SM prediction of $B\to X_{s}\gamma$ may exclude the parameter region with $m_{H^{\pm}}<800$ GeV Misiak:2020vlo , but the new HFLAV result HeavyFlavorAveragingGroupHFLAV:2024ctg would relax the bound drastically. We plot our predictions in both heavy and light charged Higgs cases. There is no big difference between in the Type-II 2HDM and in the Type-Y 2HDM, when charged Higgs is heavy. If the charged Higgs is light, the signal strength, $\mu_{\tau\tau}$, tends to be larger in the Type-Y than that in the Type-II. In our model, there are extra quarks which can contribute to $B\to X_{s}\gamma$. The study of flavor physics involving extra quarks, however, is not simple, since other flavor observables are usually correlated to $B\to X_{s}\gamma$. This is also our future work. Finally, let us comment on how to distinguish our models with the others in the previous works. As we see in Table 5, $Z^{\prime}$ mass, that is ${\cal O}(g_{X}v_{\Phi})$, is below 1 TeV in our model. There are several possibilities of $U(1)_{H}$ charge assignments and matter fields as discussed in Ref. Ko:2012hd , but in any case, the resonance search below 1 TeV play an important role in testing our models. As studied in Ref Ko:2015fxa , there is a good DM candidate in the Type-II 2HDM with gauged $U(1)_{H}$ symmetry. Our analysis in this paper does not include the DM, but this prediction is also important to test and distinguish our model with the others.